OpenStudy (anonymous):

Please help me I will do anything I have been stuck on this problem for honestly 3 weeks. A local grocery store has agreed to sell your homemade bread. You will use the following information along with some ideas from Chapter 3 to decide how many loaves should be manufactured each week and what price should be charged. After tracking weekly sales at several different prices, you get the following data: Loaves Sold, x Price, p 355/$1.50 320/$2.00 265/$2.50 235/$3.00 180/$3.50 125/$4.00 In order to increase manufacturing capacity, you’ve taken out a loan to buy an industrial sized oven for $4000. The new oven will allow you to make a maximum of about 400 loaves of bread per week. The loan is to be paid back monthly over two years at an annual interest rate of 10% compounded monthly. The monthly payments are $203.40. (You can check these numbers after section 5.7.) The ingredients for two loaves of bread are given in the table below. The $1.182 is the cost of the ingredients for a single loaf of bread. Demand Equation. Make a scatter plot of the six data points (using the number sold as the x-coordinate.) Does the relationship appear to be linear? Use regression analysis to find the line of best fit. This line will be your demand equation. How strong is the correlation? Revenue Function. Find R(x), the weekly revenue as a function of loaves sold, x. (Note that R(x) is an equation not a single value.) Cost Function. Find C(x), the weekly cost for producing x loaves of bread. Be sure to include both the cost of the oven and the ingredients. What is the domain of the cost function? Profit Function. Find P(x), the weekly profit for producing and selling x loaves of bread. (Hint: profit = revenue – cost.) Maximum Revenue. Find the number of loaves that should be sold in order to maximize revenue. What is the maximum revenue? What price should be charged in order to maximize revenue? Maximum Profit. Find the number of loaves that should be produced and sold in order to maximize the profit. What is the maximum profit? What price should be used to maximize profit? Conclusion. How many loaves of bread will you produce each week and how much will you charge for each loaf? Why?

OpenStudy (anonymous):

this looks like a pain

OpenStudy (anonymous):

haha yes i have been stumped for a week, im just hoping that someone can help me out some.

OpenStudy (anonymous):

here is a picture they wanted of the points plotted http://www.wolframalpha.com/input/?i=%28355,1.50%29+%28320,2.00%29+%28265,2.50%29+%28235,3.00%29+%28180,3.50%29+%28125,4.00%29

OpenStudy (anonymous):

Demand Equation: y=-0.01x+5.4287, this is what i got for the demand equation

OpenStudy (anonymous):

\[y=5.42873-0.0108597 x \]

OpenStudy (anonymous):

probably more decimals than you need

OpenStudy (anonymous):

so for the revenue function would i use the demand equation to find that?

OpenStudy (anonymous):

i actually have no idea it is the number of loaves sold times the price per loaf

OpenStudy (anonymous):

a quick google search provides the answer

OpenStudy (anonymous):

$3.00 is the best price point but i need to also find the max revenue and max profit for loaves of bread and that could be any number from 1-400

OpenStudy (anonymous):

can you post the link?

OpenStudy (anonymous):

the demand function is what you just got, \[y=5.42873-0.0108597 x\] hold on i got confused, let me send you an example

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

y in our case was the demand, x was the price so the revenue funtion is the demand times the price, i.e. \(xy=5.42873x-0.0108597 x^2\)

OpenStudy (anonymous):

that is \[R(x)=5.429x-0.0109x^2\] here i rounded a bit

OpenStudy (anonymous):

or you can round like you did if you prefer in any case the revenue is that times x

OpenStudy (anonymous):

how we doing so far?

OpenStudy (anonymous):

good i think that the cost function is C(x)=50.85+1.182*x

OpenStudy (anonymous):

yeah looks good, divided the monthly cost by 4

OpenStudy (anonymous):

yes that is what i did then for the profit function P(x)=5.429x-50.85+1.182*x right?

OpenStudy (anonymous):

profit is revenue minus cost right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

something screwy cause revenue had an \(x^2\) term in it which got lost somewhere

OpenStudy (anonymous):

\[5.429x-0.0109x^2-(50.85+1.182x )\]

OpenStudy (anonymous):

you will of course want to multiply out and combine like terms

OpenStudy (anonymous):

that is for the profit right because it was literally just the revenue equation- the profit equation.

OpenStudy (anonymous):

that is profit since it is "revenue minus cost"

OpenStudy (anonymous):

i used wolfram to deal the the decimals and got this \[P(x)=-0.0109 x^2+4.247 x-50.85\]

OpenStudy (anonymous):

All right and then to find the Maximum rev and profit dont you just make that equation equal to 0?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

that will tell you when you get zero profit or zero revenue you certainly don't want that !!

OpenStudy (anonymous):

which one do you want to find, the max of the revenue function or that max of the profit function (or both)?

OpenStudy (anonymous):

well i have to find both of them for this project but we can do max revenue first.

OpenStudy (anonymous):

ok the max of the revenue function is the second coordinate if the vertex you find it by finding the first coordinate of the vertex, and plugging it it to get the second coordinate in this case the first coordinate represents the number of loaves and the second represents the max revenue

OpenStudy (anonymous):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\)

OpenStudy (anonymous):

\[R(x)=5.429x-0.0109x^2\] \[a=-0.0109,b=5.429\] use a calculator to compute \[\frac{5.429}{2\times 0.0109}\]

OpenStudy (anonymous):

Ok so i got 249.04 for that

OpenStudy (anonymous):

yeah i got that too, i would go with 249 as it is loaves of bread

OpenStudy (anonymous):

then to get the revenue, plug it in

OpenStudy (anonymous):

i would cheat

OpenStudy (anonymous):

so for revenue it would be 5.429(249)-0.0109x^2 right

OpenStudy (anonymous):

scroll down to where it says "global max" then click on "approximate form" you will see that it is about 696 when x is about 249

OpenStudy (anonymous):

no you have to replace both x's by 249

OpenStudy (anonymous):

\[5.429(249)-0.0109\times 249^2 \]

OpenStudy (anonymous):

oops forgot to send the link http://www.wolframalpha.com/input/?i=5.429x-0.0109x^2

OpenStudy (anonymous):

So i got 676 for the maximum rev but you said 696 earlier did i do something wrong?

OpenStudy (anonymous):

maybe i misread it

OpenStudy (anonymous):

yes, you are right and i am wrong 676

OpenStudy (anonymous):

oh ok and then to find the Maximum profit what should i do?

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

but this time for the profit function instead of the revenue function compute \(-\frac{b}{2a}\) for \[P(x)=-0.0109 x^2+4.247 x-50.85\]

OpenStudy (anonymous):

would that be -50.85/2(-0.0109)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(b=4.247\) not \(-50.85\)

OpenStudy (anonymous):

and there is a minus sign in the formula, so the minus signs will go and the answer will be positive (good thing too)

OpenStudy (anonymous):

\[4.247\div(2\times 0.0109)\]

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

alright for that i got 194.82

OpenStudy (anonymous):

me too might want to round to 195 since they are whole loaves

OpenStudy (anonymous):

and now i have to go, hope you are goo from there at least it did not take us three weeks to do this!

OpenStudy (anonymous):

more than an hour, but not three weeks!

OpenStudy (anonymous):

haha thank you so much you don't even realize how much trouble I have had with this you are a life saver.

OpenStudy (anonymous):

yw